OpenStudy (student_123456789_):

Finding arc length by calculating definite integral of (1+[f'(x)]^2)^(1/2)dx for x=-16 to x=16 if f(x) = 19.997/2(e^(x/19.997)+e^(-x/19.997))-19.997 showing steps

zepdrix (zepdrix):

Hey Student! Welcome to OpenStudy! Hmm your function is a little difficult to read in text, is this what it should look like?\[\large\rm f(x)=\frac{19.997}{2}\left(e^{x/19.997}+e^{-x/19.997}\right)-19.997\]

OpenStudy (student_123456789_):

yes..sorry

zepdrix (zepdrix):

Do you understand how to find your derivative?\[\large\rm f'(x)=?\]

OpenStudy (student_123456789_):

Yes ...I think in its simplest form it is f'(x) = \[(1/2)(e^(x/19.997)-e^(-x/19.997))\] I understand differentiation and integration and when I integrate using this f'(x) in the integral function on my calculator I get the answer which should be close to 36. But i need to show it algebraically and cant get it...,

zepdrix (zepdrix):

Just a note for using the equation tool/ LaTeX: Use curly braces if you want to put more than one character into an exponent. Example: e^{4x}=\(e^{4x}\)

zepdrix (zepdrix):

Algebraically, we can apply an identity,\[\large\rm \sinh(x)=\frac{1}{2}\left(e^x-e^{-x}\right)\]

zepdrix (zepdrix):

\[\large\rm f'=\frac12\left(e^{x/19.997}-e^{-x/19.997}\right)\]Do you see how we can apply this identity to our derivative? :)

OpenStudy (student_123456789_):

i see the connection but I haven't learnt that identity so I don't know how to work with it....

zepdrix (zepdrix):

Hmm I'm not sure if this will work out nicely otherwise >.< Thinking...

zepdrix (zepdrix):

Ya I'm not really seeing another way to approach this. \[\large\rm f'=\sinh\left(\frac{x}{19.997}\right)\]If this approach is too confusing, new, we can keep trying stuff maybe... Lemme know what you're thinking XD

OpenStudy (student_123456789_):

can I just check that if I were to integrate it I would do ((1+(f'(x)^2)^(3/2) and then divide by [(3/2) x the derivative of what was inside the root] ? just want to make sure I am following the correct rule for integration

OpenStudy (student_123456789_):

The thing is is that I am doing a school task that other students have to do and none of us have learnt that identity so I am just assuming there must be a way to do it without it? sorry to be difficult... thanks for helping

zepdrix (zepdrix):

No that wouldn't quite work. Take a simple example like:\[\large\rm \int\limits \sqrt{1+x^2}~dx\quad\ne\quad \frac{2}{3(2x)}(1+x^2)^{3/2}\]It's not even close to that actually. You would need to apply a trig sub,\[\large\rm \int\limits \sqrt{1+(f')^2}~dx\]\[\large\rm f'=\tan \theta\]\[\large\rm \int\limits \sqrt{1+\tan^2\theta}~dx\]The problem with that approach is... replacing our differential dx is going to be rather hard. Maybe there is a clever trick to it that I'm just not thinking of.

zepdrix (zepdrix):

\[\large\rm f'=\frac12\left(e^{x/19.997}-e^{-x/19.997}\right)\]Squaring,\[\large\rm (f')^2=\frac14\left(e^{x/19.997}-e^{-x/19.997}\right)^2\]\[\large\rm (f')^2=\frac14\left(e^{2x/19.997}-2+e^{-2x/19.997}\right)\]Adding 1 and square rooting,\[\large\rm \sqrt{1+(f')^2}=\sqrt{1+\frac14\left(e^{2x/19.997}-2+e^{-2x/19.997}\right)}\]Maybe common denominator,\[\large\rm \sqrt{1+(f')^2}=\sqrt{\frac44+\frac14\left(e^{2x/19.997}-2+e^{-2x/19.997}\right)}\]Which we can write as,\[\large\rm \sqrt{1+(f')^2}=\frac{1}{2}\sqrt{4+\left(e^{2x/19.997}-2+e^{-2x/19.997}\right)}\]OH OH OH yes I'm seeing it!

zepdrix (zepdrix):

I'll stop there a sec, I don't wanna ruin any of the fun if you see it. Read that a sec and lemme know if any of it's confusing.

OpenStudy (student_123456789_):

sorry so I understand what you have done but I don't know what to do next....

zepdrix (zepdrix):

Not jumping out at you? That's ok! It's a tricky little trick they're pulling here. So let's combine the 4 and the -2,\[\large\rm =\frac{1}{2}\sqrt{e^{2x/19.997}+2+e^{-2x/19.997}}\]

OpenStudy (student_123456789_):

yep I get that then can I try and get it all squared so I can remove the root...will that work?

zepdrix (zepdrix):

Good good good. When we had the -2 in the middle, it was a perfect square. So the +2 in the middle probably gives us some other perfect square!

OpenStudy (student_123456789_):

oh ok so its just the perfect square but with a plus in the middle and then the root cancels with the square and then its just 1/2 (e^2x/19... +e^-2x/19..) and then just integrate that?

zepdrix (zepdrix):

\[\large\rm e^{2x/19.997}+2+e^{-2x/19.997}\quad=\quad \left(e^{x/19.997}+e^{-x/19.997}\right)^2\]Mmm yes very good. Then it's as simple as integrating a couple of exponentials, ya?

OpenStudy (student_123456789_):

yes got 35.525 as length! thank-you!!!

OpenStudy (student_123456789_):

also I'm new to this - am I supposed to like reward you with something...? I just dont really know this site works

zepdrix (zepdrix):

You can click the "Best Response" button by my name if you found my responses helpful. It doesn't really matter though :) Ooo you got it? Noiceeee.

OpenStudy (student_123456789_):

ok thanks again!

zepdrix (zepdrix):

np